### v4

@@ 2.4   Interpreting the parallel trends assumption in the outcome
Thus far, we have formalized the target parameter and the identifying assumptions in $2\times 2$ DID-IV designs. Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand to identify the effects of the instrument on the treatment and the outcome. The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs. To do so, we first uncover the selection mechanism behind the canonical DID-IV designs. Specifically, we investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups. Then, we show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment. We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path is $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
Recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods. In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods.
Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers. The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods. In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument). However, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument; that is, their treatment choices are affected by time in the absence of a policy shock.
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1). We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
@@
We now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in the exposed and unexposed groups, respectively.
@@
Note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$. At the same time, $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
@@
$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$
$\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$
$\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument, respectively, for the time compliers in an unexposed group.
Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$). We can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit. However, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservable in an exposed group ($e=1$) because we cannot specify the time group of unit $i$ in that group. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we cannot distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},$
$\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (the stationary conditions hold), these quantities are equal to zero. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
@@
The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the expectation of the time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively. The second term represents the selection gain in each time group in group $E_{i}=e$, respectively. The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
@@
See Appendix A. ∎
@@
Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ is equal to a weighted average of the indirect effect of time on the outcome through treatment for each time group $g^{t}$ in group $e$. The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$. In this paper, we call $\Delta_{e}$ an expected time gain in group $e$ hereafter.
@@
Based on the above discussions, we conclude that the parallel trends assumption in the outcome (Assumption 6), that is, $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups; the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
@@
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period. In practice, if we have pre-exposed period data, we can indirectly check the validity of Assumption 6 by comparing the evolution of the outcome between exposed and unexposed groups. The formal discussion for the pre-trend test is provided in Subsection 5.3.
2.5   Relationship between DID and DID-IV
In this subsection, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with noncompliance of the assigned treatment.
$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (1)
These restrictions imply that we have Tables 3-4 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 3 - 4:
Table 3: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
Table 4: Treatment group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
$=LATET.$
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Tables 3- 4, we can rewrite the parallel trends assumption in the untreated outcome as:
$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
$\iff$ $\Delta_{0}=\Delta_{1},$
Overall, this subsection clarifies that DID-IV is a natural generalization of DID, and the differences between the two designs arise because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$. In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
@@ 2.6   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture some causal effects under the same setting as in this article. However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID. The detailed discussion is given in Appendix E, and this section provides the overview of that discussion.
The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see Subsection 2.2). Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In Appendix E, we first show that the identifying assumptions in Fuzzy DID designs ex-ante impose the strong restrictions on the treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 5-6444Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6., which are in line with Tables 1-2. From these tables, one can see that dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group. Note that in our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption. The difference between dCDH’s restrictions and ours is critical for empirical application; in contrast to the monotonicity assumption, which excludes the defiers in both groups, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application. This is because these restrictions ex ante exclude the units who change the treatment adoption in the absence of a policy shock (instrument).
@@
Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers. The other causal parameter measures the treatment effects among the time compliers $CM^{T}$. The time compliers switch from non-treatment to treatment in period $1$ in the absence of the policy shock (instrument); that is, they are affected by time but not affected by the instrument. This decomposition result for the SLATET has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET. Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided. During this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups for the Wald-DID estimand to identify the SLATET.
@@
First, we show that dCDH’s argument regarding the use of the Wald-DID estimand is based on their reliance on the parallel trends assumption in DID designs. Recall that in our DID-IV design, we did not make the parallel trends assumption in the untreated outcome, but made the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. In the proof, we also demonstrate that imposing the parallel trends assumption in DID designs yields the strong restrictions on the treatment adoption behavior across units under Fuzzy DID designs.
@@
Next, after confirming that dCDH’s argument is based on the parallel trends assumption in DID designs, we argue that this assumption is unsuitable for DID-IV settings for three reasons. First, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires the data to contain parents with zero education attainment, which is unrealistic in practice. Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00. We show that dCDH’s placebo test is incomplete for assessing the validity of this assumption. Surprisingly, we also show that their placebo test coincides with our pretrends test with a slight modification, introduced in section 5.3, to assess the plausibility of Assumption 5 and Assumption 6.

### v5

@@ 2.4   Interpreting the parallel trends assumption in the outcome
In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs. It is less clear than that in DID designs because this assumption does not depend on the untreated outcome, but depends on the unexposed outcome.
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$ that represents the treatment adoption process during the two periods if the instrument path is $z=(0,0)$. Recall that in DID-IV designs, the units are allowed to take the treatment both period 00 and $1$ without instrument. In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods. Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
@@
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 3 (monotonicity assumption).
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. [^9] In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1). We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
We now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively.
@@
Note that $\Delta_{0}$ is estimable by the observed change in the outcome in an unexposed group during the two periods. By contrast, $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
@@
$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$ $\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$ $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
We can interpret each $\Delta_{g^{t},e}$ as the effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},\hskip 5.69054pt\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
@@
The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the average time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively. The second term represents the selection gain in each time group in group $E_{i}=e$, respectively. The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
@@
This directly follows from the Law of Iterated Expectation:
$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
Clearly, the weight $w_{g^{t},e}$ sums to one by construction. ∎
@@
Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ can be written as the weighted average of the expectation of the potential outcomes, and can be interpreted as measuring the convex average of the time effect for each time group $g^{t}$ in group $e$. The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$. Hereafter, we call $\Delta_{e}$ an expected time gain in group $e$.
@@
From Theorem 2, we can interpret that the parallel trends assumption in the outcome (Assumption 6), $\Delta_{0}=\Delta_{1}$, requires the same expected time gain between exposed and unexposed groups; that is, the effects of time on the outcome through treatment are the same between the two groups.
@@
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
Finally, we provide an empirical example to illustrate the interpretation of each notation and Assumption 6 in practice.
Example.
Suppose that the treatment $D_{i,t}\in\{0,1\}$ represents participation in a job training program, the outcome $Y_{i,t}$ represents real earnings, and the instrument $Z_{i,1}\in\{0,1\}$ represents a policy shock that encourages people to receive the treatment in the second period. [^10] Assume that the treatment is non-staggered and the treatment and the instrument are not randomized across units. In this setting, the instrumental variable strategy nor the canonical DID designs are directly applicable. In this setting, the time compliers $CM^{T}$ are the units who start adopting the program in the second period in the absence of the policy shock, and $w_{CM^{T},0}$ measures its share in an unexposed group. $\Delta_{CM^{T},0}$ represents the average time trends of real earnings among the time compliers in an unexposed group, and can be decomposed into the time effects on real earnings without policy shock, and the causal effects of program participation on real earnings in the second period, as their treatment adoption behavior is affected by time. Assumption 6 requires that the average time trends of real earnings are the same without policy shock between exposed and unexposed groups. This assumption can be interpreted as the restriction on the effect of time on real earnings through program participation between the two groups.
@@ 2.5   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting as in this article. [^11] The Wald-DID estimand considered in dCDH corresponds to ours in this article. In their empirical application, dCDH revisit Duflo (2001), in which the group variable $E_{i}$ ($G_{i}$ in dCDH) takes one if unit $i$ belongs to the region that are highly exposed to the policy shock. However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID in an overview of the detailed discussion provided in online Appendix Section E.
The main difference between dCDH and this article lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.121212de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. Therefore, we rewrite their target parameter by the notation used in this article.
Definition.
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
where the switcher (SW) is the units who become treated in period $1$.
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). Because the target parameter is different from each other, the identifying assumptions in Fuzzy DID designs are also different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In online Appendix Section E, we first show that the identifying assumptions in Fuzzy DID designs ex ante impose strong restrictions on treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups. Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 3-4, which are in line with Tables 1-2.131313Because dCDH consider repeated cross section settings, we can observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.
Comparing Tables 1-2 with Tables 3-4, one can figure out the difference between dCDH’s restrictions and ours. In our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption. By contrast, dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
@@
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by dCDH’s identifying assumptions.
@@
Next, under Tables 3-4, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers. The other causal parameter measures the treatment effects among the time compliers $CM^{T}$. [^14] Intuitively, we cannot identify this parameter in DID-IV settings because the untreated potential outcomes in period $t=1$ among the time compliers are not observed in both groups (see Tables 3-4). This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher, as this parameter is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by the instrument.
Finally, we revisit dCDH’s argument regarding the use of the Wald-DID estimand, and show that this problem is less serious than one might expect, and does not arise once a researcher imposes Assumption 6 instead of the parallel trends assumption in the untreated outcome.
@@
First, we show that their argument is based on the parallel trends assumption in DID designs. Recall that in our DID-IV design, we do not make the parallel trends assumption in the untreated outcome, but make the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
@@
Next, we argue that the different type of the parallel trends assumption, Assumption 6, is suitable for DID-IV settings, and the parallel trends assumption in the untreated outcome would be difficult to justify in practice for three reasons. First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Duflo (2001), this assumption requires the data to contain units with zero education attainment, which is unrealistic in practice. Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.